summerh0819:

What is (f-g)(x) f(x)= x^3-2x^2+12x-6 g(x)= 4x^2-6x+4 Enter your answer in standard form in the box. (f-g)(x)=

AZ:

Another way to write that is (f - g)(x) = f(x) - g(x)

AZ:

So (x^3-2x^2+12x-6) - (4x^2-6x+4) First we have to distribute the negative inside the parenthesis, can you do that?

summerh0819:

I'm so lost lmao

snowflake0531:

better way to ask do you know how to add/subtract polynomials?

summerh0819:

No not really

snowflake0531:

oop

summerh0819:

My brain is legit fried rn I've been at work all day and coming home to do schoolwork is not my favorite thing to do especially MATH.

snowflake0531:

well, then thank @AZ for being here lol

AZ:

let me show you an example with some simpler numbers if we have two functions, one is called f and another is called g f(x) = x + 2 g(x) = 2x + 3 and they want us to find (f-g)(x) as I said earlier (f-g)(x) is the same thing as f(x) - g(x) so that would be (x+2) - (2x+3) we have to distribute the minus inside which means we flip the signs inside and we'd get x + 2 - 2x - 3 and then we can simplify it

summerh0819:

Thank yall both for helping me

AZ:

ah, yes doing math after a long day is no fun I'm not sure if you've read my example so far, but it's fine if you didn't let's take it step by step

summerh0819:

Okay thank you, i'm reading it now

AZ:

So we said we had to do \(\color{red}{f(x)} - \color{orange}{g(x)}\) and we took that huge chunk of letters and numbers and put them in their place and got \(\color{red}{(x^3-2x^2+12x-6)} - \color{orange}{(4x^2-6x+4)}\) but let's look at it one part at a time I added colors so you could see which part is f(x) and which one is g(x) and not get confused because there's so many letters and numbers everywhere

summerh0819:

Thanks lol.

AZ:

the first step we have to do is- distribute the negative sign inside of the parenthesis so let's only look at the right half that's in orange \(- (4x^2 - 6x + 4)\) do you know how to multiply that negative sign inside? as a hint using some other letters -(a + b - c) = -a - b + c do you see how the signs in front of each variable (or letter) changes?

summerh0819:

yeah

AZ:

good! so can you similarly do that for −(4x^2 − 6x + 4)

summerh0819:

ok

AZ:

 4x^2 is positive so it's going to become -4x^2 when we bring that - inside the parenthesis what's the opposite of the minus sign? is it a + sign? and for the third one- what's the opposite of the plus sign